OpenStudy (anonymous):

Find the measures of two angles, one positive and one negative, that are coterminal with pi/5 .

OpenStudy (anonymous):

you can add \(2\pi\) to \(\frac{\pi}{5}\) to find a positive one, then subtract \(2\pi\) from \(\frac{\pi}{5}\) to find a negative one

OpenStudy (anonymous):

so 6pi/5 and -4pi/5?

OpenStudy (anonymous):

btw don't be annoyed by the \(\pi\) when adding the arithmetic for \(\frac{\pi}{5}+2\pi\) is identical to the arithmetic for \(\frac{1}{5}+2\)

OpenStudy (anonymous):

no, \(\frac{1}{5}+1=\frac{6}{5}\)

OpenStudy (anonymous):

11pi/5 and -pi/5

OpenStudy (anonymous):

first one is right, because \(2+\frac{1}{5}=\frac{11}{5}\)

OpenStudy (anonymous):

but the second one is wrong, becaues \(\frac{1}{5}-2\neq -\frac{1}{5}\)

OpenStudy (anonymous):

it could also be -9pi/5

OpenStudy (anonymous):

yes, that is the correct answer

OpenStudy (anonymous):

\(\frac{1}{5}-2=-\frac{9}{5}\) and so \[\frac{\pi}{5}-2\pi=-\frac{9\pi}{2}\]

OpenStudy (anonymous):

Find the measures of two angles, one positive and one negative, that are coterminal with pi/10 .

OpenStudy (anonymous):

same idea exactly

OpenStudy (anonymous):

\(\frac{1}{10}+2\) and \(\frac{1}{10}-2\) then stick a \(\pi\) in your answer

OpenStudy (anonymous):

21pi/10 and -19pi/10